# Automated Bug Generation in the era of Large Language Models

### v1

@@ 21       $\mathit{i}\leftarrow i+1$;
4.3. Bug Generator
Bug generation involves modifying, adding, or deleting code segments from an original method to change the expected behavior of the program. The Bug Generator module of BugFarm takes a method and its corresponding set of $LAS$ identified in the previous step as inputs, and crafts LLM prompts to generate $N$ buggy versions for the method as outputs. Our intuition for making BugFarm configurable is that our bugs, very likely, will be used as training/fine-tuning data for bug-related tasks. So, generating multiple buggy versions of a single method would be helpful for a model to distinguish between buggy and non-buggy code easier. The total number of bugs, however, also depends on the size of the method and threshold value $k$. For example, $k=10\%$ for methods with less than $10$ statements returns one statement as LAS, and changing that statement in $N$ unique ways may be infeasible.
BugFarm’s prompts consist of three parts. The first part is the natural language instruction asking LLM to generate the bugs. The second part provides contextual information about where to inject the bug, i.e., only to consider the least attended statements, $LAS$, for making bug-inducing changes. Finally, we include the entire method, including both signature and the method body, in the prompt. Such prompts are LLM-agnostic, i.e., they can be used with various existing LLMs, such as LLaMa (Meta, 2023), PaLM (Google, 2023b), Copilot (GitHub, 2023), Alpaca (Taori et al., 2023), and ChatGPT (AI, 2023). The only consideration is to check if the size of the prompt matches the context window of the LLM.
@@ 4.3. Bug Generator
6 foreach $Response\in Responses$ do
@@ 15                   $break$;
After the LLM’s response to the prompts, Bug Generator component validates (1) if they are syntactically correct and (2) if the changes do not impact the attention of the model. The responses that do not pass these two checks will be discarded. Ideally, we want the produced bugs to be syntactically correct. To that end, our prompts also include instructions to produce syntactically correct code. However, our experiments showed that despite asking an LLM to produce syntactically correct code, specifically for longer methods, it generates responses that are not parsable. Furthermore, our goal is to generate bugs that have a very close neural representation with the original code. So, we assess the model’s attention to the generated bugs and only accept them if the changes are still among the least attended statements.
Algorithm 2 demonstrates BugFarm’s bug generation and validation approach. It starts by initializing the output variable $Bugs$ (Line 1) and getting the indices of LASs in $method$ (Line 2). Next, it adds an index to each statement in $method$ (Line 3). This will help us to refer to LASs in the prompt by number instead of including the whole statements, resulting in a reduction of the prompt size. This is specifically important for longer methods or statements, as the context window of LLMs is limited (Nguyen and Wong, 2023; Xu et al., 2022). Next, we will craft the prompt with the required context, i.e., indexed method and bug injection location, and send the prompt to an LLM (Lines 4-5)  [^2] The natural language part of our prompts is more complex than the example here. The readers can refer to our artifact to see the exact prompts (authors, 2023). Finally, once the LLM responds, we check if the generated bugs are syntactically correct (Lines 7-8) and if the changed statements are among the least attended statements by the model (Lines 9-14) to add them to the acceptable set of bugs (Lines 15-16).
5. Evaluation
To evaluate the effectiveness of BugFarm, we compare it with two state-of-the-art automated bug generation approaches, namely LEAM and $\mu$BERT, investigating the following research questions:
@@ 1. RQ1:
    Characteristics of the Generated Bugs. Can BugFarm successfully inject bugs into any arbitrary code? What are the characteristics of the generated bugs by each approach? To what extent do our bugs overlap with those generated by other approaches?
@@ 3. RQ3:
    Effectiveness in Generating Hard-to-Repair Bugs. To what extent can a learning-based Automated Program Repair (APR) technique repairs BugFarm bugs compared to that generated by alternative approaches?
@@ 4. RQ4:
    Necessity of Prompt Engineering. What are the benefits of prompt engineering in BugFarm?
  5. RQ5:
@@
5.1. Experimental Setup and Data Availability
@@
Alternative Approaches. We compare BugFarm with two most recent mutant generation techniques, $\mu$BERT (Khanfir et al., 2023) and LEAM (Tian et al., 2022). To inject mutants, $\mu$BERT selects AST nodes representative of program behavior—literals, identifiers, expressions, assignments, object fields, method calls, array access, and types. Then, it replaces the tokens in selected AST nodes with the spacial token <mask> and uses CodeBERT to predict the masked token. The intuition is that if CodeBERT predicts a token different from the original one, the transformation introduces a bug. LEAM is a deep learning-based technique that learns to mutate code from large examples of real-world bugs. To that end, they represent code as a sequence of AST nodes and learn to apply eight grammar rules to select and modify the code. Both $\mu$BERT and LEAM claim to generate better bugs (mutants confirmed by test execution) compared to classic approaches, namely PIT (Coles et al., 2016) and Major (Just, 2014). Therefore, we did not include classic techniques in our evaluation.
@@
For BugFarm, we used the threshold value $k=10\%$, mainly because alternative approaches do not change more than a handful of statements in a given code (more details in section 5.2). Our experimental results in the rest of this section confirm that even with such a low threshold, we still surpass other techniques. We expect a higher threshold will make our bugs more complex compared to other approaches, improving margins. We also configured BugFarm to generate at most three mutants per method ($N=3$ in Algorithm 2) due to budget allowance. The current implementation of BugFarm’s Bug Generator component uses GPT-3.5-turbo as an LLM due to its superiority over a large number of models (Bubeck et al., 2023; Zhang et al., 2023; OpenAI, 2023; Pan et al., 2023) and cost-effectiveness. Using a more advanced LLM such as GPT-4 likely yields better results, not worse.
Bug Validation. Alternative approaches we compare against generate mutant, which may not necessarily change the outcome of test execution. Worse, the mutants may not even be compilable. For BugFarm, although we ask the LLM to generate syntactically correct bugs, it is still not guaranteed that the generated code executes and changes the outcome of test execution. This entails implementing a validation process to confirm if the generated mutants/bugs are compilable and can change the outcome of test execution to consider them bugs. Our validation procedure follows these steps: We first run existing test suites on the original code and select the green tests. Next, we take all the generated mutants/bugs by different techniques, compile them, and run the previously green tests against them (each mutant will be tested in isolation). We discard those that do not compile or result in no failure after test execution. All the experiments answering RQ1-RQ5 are done with confirmed bugs.
@@
Bug Prediction Models. When selecting bug prediction models, we had to consider the following criteria: (1) the current implementation of BugFarm’s Attention Analyzer component works on transformer-based models, including customized trained models and those fine-tuned on pre-trained code-language models; (2) we require the availability of models to perform the attention analysis, so closed-source models such as Codex (Chen et al., 2021) are off the table; (3) to compare against other bug prediction approaches, we require models with either pre-trained models or required artifacts for training available. We could not find any custom bug prediction model publicly or per request available with such characteristics. Consequently, we chose three pre-trained code-language models that are widely used by the research community, namely CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), and NatGen (Chakraborty et al., 2022) [^3] Fine-tuning larger models requires non-trivial computing resources. Our experiments will show that models superior in other code-related tasks (e.g., CodeT5 over CodeBERT) show the significance of BugFarm better. We expect this to hold for larger models as well.. We fine-tuned these models using the real world and synthetic bugs, i.e., those generated by LEAM and $\mu$BERT, and compared their effectiveness in predicting our bugs.
@@
CodeBERT is an encoder-only transformer model based on BERT (Devlin et al., 2018) architecture, which is trained on $2.1M$ bi-modal (natural language and code pairs) and $6.4M$ uni-modal (code only) data from CodeSearchNet (Husain et al., 2019) dataset. The main learning objectives in CodeBERT are Masked Language Modeling (MLM)—the model learns to predict the tokens replaced by a special mask token—and Replaced Token Detection (RTD)—the model learns to detect which token does not belong to the original data. CodeT5 is an encoder-decoder transformer model based on T5 (Raffel et al., 2020) architecture. CodeT5 is trained on $8.35M$ functions from various programming languages provided by CodeSearchNet (Husain et al., 2019) and BigQuery (Big, 2023) dataset, and its training objectives include masked span and masked identifier prediction. Such objectives enable the model to understand code semantics better than CodeBERT.
@@
NatGen is also an encoder-decoder transformer model, trained on a generative task of naturalizing source code. Specifically, NatGen starts with CodeT5—based model—parameters and continues the training with a new objective, i.e., re-constructing the original code (natural) given transformed code (de-natural). It uses $8.1M$ pairs of natural and de-natural functions from CodeSearchNet (Husain et al., 2019) and C/C#. We used the base model of CodeBERT, CodeT5, and NatGen for our experiments.
@@
Automated Program Repair Model. To evaluate the effectiveness of learning-based techniques in repairing generated bugs, we used FitRepair (Xia et al., 2023). FitRepair is a state-of-the-art APR technique that outperforms existing approaches. It leverages information retrieval and static analysis to implement domain-specific fine-tuning and prompting strategies. We used FitRepair with CodeT5-Large in a zero-shot manner to generate patches for $\mu$BERT, LEAM, and BugFarm bugs and validated if the patches thoroughly fix the bugs through test execution.
@@
Subjects. BugFarm is programming-language agnostic; none of its components depend on a specific programming language. However, we chose Java projects in our experiments for the following reasons: (1) LEAM’s pre-trained model is on Java, and they have no alternative training dataset for other programming languages; (2) we needed real-world bug datasets for RQ2, and most of the existing real-world bug datasets are in Java. For a fair comparison, we used Defects4J $V2.0$ projects as a baseline for bug generation since the alternative approaches are shown to work on them with no issues. The current version of BugFarm supports Maven projects only, so we excluded Mockito and Closure projects from the subjects. The first two columns of Table 1 show the list of our $15$ subjects and the number of methods per subject used for bug generation.
@@
Table 1. Comparing the characteristics of bugs generated by BugFarm, $\mu$BERT, and LEAM. M: # Methods, UB: # Unconfirmed Bugs, CB: # Confirmed Bugs, SI: # Statements Involved in bug generation, LD: Lines Deleted in bug generation, ED: Edit Distance, OL: Overlap with LEAM, OM: Overlap with $\mu$BERT.
Subjects M BugFarm LEAM (Tian et al., 2022) $\mu$BERT (Khanfir et al., 2023)
UB CB SI LD ED OL OM UB CB SI LD ED UB CB SI LD ED
cli 276 438 263 (60%) 2.86 0% 32.14 ¡14%,0.64¿ ¡8%,0.67¿ 3959 1375 (34.73%) 2.56 13.82% 26.2 12641 1528 (12%) 2.00 0.07% 19.38
codec 901 1586 357 (23%) 3.12 0% 32.39 ¡3%,0.67¿ ¡4%,0.85¿ 13603 2114 (16%) 2.59 7.95% 20.33 42529 9541 (22%) 2.01 0% 17.89
collections 4440 7373 2958 (40%) 2.55 0% 27.01 ¡7%,0.67¿ ¡5%,0.73¿ 56196 9356 (17%) 2.71 8.87% 21.33 132009 21367 (16%) 2.01 0% 20.02
compress 4123 6544 464 (7%) 3.04 0% 30.29 ¡5%,0.69¿ ¡2%,0.80¿ 61488 9085 (15%) 2.57 12.87% 24.56 167259 22657 (14%) 1.97 0% 23.83
csv 248 379 250 (66%) 2.66 0% 24.30 ¡12%,0.66¿ ¡1%,0.75¿ 3560 1181 (33%) 2.83 14.65% 33.60 818 106 (13%) 1.98 0% 6.81
jxpath 1672 2601 912 (35%) 3.11 0% 26.09 ¡6%,0.69¿ ¡5%,0.78¿ 26615 5645 (21%) 2.86 16.44% 26.81 90153 10678 (12%) 2.00 0% 25.88
lang 3810 6505 3734 (57%) 2.95 0% 29.76 ¡7%,0.68¿ ¡3%,0.76¿ 54101 15861 (29%) 2.66 10.01% 21.60 80200 15618 (19%) 2.00 0.02% 21.13
math 5796 8935 6104 (68%) 3.27 0% 33.12 ¡4%,0.67¿ ¡2%,0.77¿ 92243 40299 (44%) 2.70 8.56% 17.15 359023 83206 (23%) 2.01 0.01% 18.30
gson 985 1497 537 (36%) 3.19 0% 36.43 ¡8%,0.7¿ ¡2%,0.68¿ 14496 2957 (20%) 2.60 18.16% 27.89 25359 2067 (8%) 2.00 0.05% 19.84
jackson-core 2626 4494 1707 (38%) 3.85 0% 41.85 ¡3%,0.68¿ ¡2%,0.72¿ 44364 9697 (22%) 2.43 22.74% 25.74 62963 9296 (15%) 2.09 0% 21.58
jackson-db 8076 12812 3855 (30%) 3.76 0% 41.70 ¡6%,0.66¿ ¡5%,0.74¿ 115859 17789 (15%) 2.78 12.61% 28.96 254947 22300 (9%) 2.01 0% 22.38
jackson-xml 586 877 249 (28%) 3.80 0% 44.24 ¡6%,0.68¿ ¡5%,0.71¿ 9935 1541 (16%) 2.67 13.17% 31.46 16420 1225 (7%) 1.98 0.08% 14.64
jfreechart 8602 13679 3035 (22%) 2.88 0% 25.26 ¡5%,0.71¿ ¡0.6%,0.81¿ 164460 16716 (10%) 2.44 11.23% 19.64 111341 7026 (6%) 2.01 0.07% 28.20
joda-time 4279 7833 3574 (46%) 2.38 0% 20.61 ¡6%,0.66¿ ¡2%,0.74¿ 69548 13739 (20%) 2.87 7.01% 24.73 141682 28010 (20%) 2.00 0% 21.21
jsoup 1642 2694 1335 (50%) 2.79 0% 28.47 ¡9%,0.66¿ ¡4.5%,0.7¿ 22632 5311 (23%) 2.52 14.54% 23.54 30420 4002 (13%) 2.24 0.05% 21.07
Total 48062 78247 29334 (37%) - - - - - 753059 152666 (20%) - - - 1527764 238627 (16%) - - -
Average 3204 5216 1956 (-) 3.08 0% 31.58 ¡7%,0.67¿ ¡3.49%,0.75¿ 50204 10178 (-) 2.65 13% 24.90 101851 15908 (-) 2.00 0.02% 20.14
@@
Evaluation Metrics. To compare the performance of bug prediction models, we use accuracy, precision, recall, F1 score, False Positive Rate (FPR), and False Negative Rate (FNR) as our metrics. To evaluate the APR results, we measure the repair success rate, i.e., the number of bugs the technique successfully patches. We define TP, TN, FP, and FN as below:
  • True Positive (TP). The code is buggy; the model predicts it as buggy.
  • True Negative (TN). The code is not buggy; the model predicts it as non-buggy.
  • False Positive (FP). The code is not buggy; the model predicts it as buggy.
  • False Negative (FN). The code is buggy; the model predicts it as non-buggy.
Data Availability. The implementation of BugFarm and all the artifacts required for reproducing the results presented in this paper are publicly available (authors, 2023).
5.2. RQ1: Quality of the Generated Bugs
To answer this research question, we compared the characteristics of the bugs generated by BugFarm with alternative approaches using the metrics listed below. The metrics are computed for all the methods and averaged for each project. For BugFarm, we have also averaged the numbers across all the models (CodeBERT, CodeT5, and NatGen) [^4] BugFarm data per base model is publicly available in the artifact website..
  • # Unconfirmed bugs: This metric, computed under columns UB in Table 1, shows the total number of unconfirmed bugs/mutants generated for each project. All the approaches successfully generate mutants/bugs for all the methods in the subject projects. $\mu$BERT generates $1.5$ and $8$ times more bugs compared to LEAM and BugFarm, respectively.
  • # Confirmed bugs: This metric, shown under columns CB in Table 1, measures the number of confirmed bugs per each project generated by different approaches. The numbers inside the parenthesis indicate the percentages of confirmed bugs compared to unconfirmed bugs. A higher percentage implicitly implies the superiority of the technique in bug generation, i.e., the code transformations are more likely to be bugs. From these results, we see that 37% of the generated bugs by BugFarm are real bugs, compared to 20% and 16% for LEAM and $\mu$BERT, respectively. Except for one project (Commons Compress), the #CB/#UB ratio for BugFarm is higher, with a significant margin. The rest of the metrics are computed for confirmed bugs, which we refer to as bugs for simplicity.

### v2

@@ 28       $\mathit{i}\leftarrow i+1$;
With the least attended tokens extracted, the algorithm can identify the least attended statements, $LAS$. To that end, it weighs each statement by a score (Lines 11-14), which is the ratio of the number of least attended tokens in that statement, normalized by its length. The key idea here is that a statement with the highest overlap between least attended tokens should achieve a lower score and, thus, be considered the least attended statement. Without normalizing the statement length, longer statements will be penalized, i.e., BugFarm never selects them to mutate. Finally, the statements are sorted in ascending order based on their scores, and the least $k\%$ attended statements (we take the ceiling in case $k\%$ of total statements is less than one) will be returned as $LAS$ (Lines 15-20).
IV-C Bug Generator
Bug generation involves modifying, adding, or deleting code segments from an original method to change the expected behavior of the program. The Bug Generator module of BugFarm takes a method and its corresponding set of $LAS$ identified in the previous step as inputs and crafts LLM prompts to generate $N$ buggy versions for the method as outputs. Our intuition for making BugFarm configurable is that our bugs will likely be used as training/fine-tuning data for bug-related tasks. So, generating multiple buggy versions of a single method would be helpful for a model to distinguish between buggy and non-buggy code more easily. The total number of bugs, however, also depends on the size of the method and threshold value $k$. For example, $k=10\%$ for methods with less than $10$ statements returns one statement as LAS, and changing that statement in $N$ unique ways may be infeasible.
1
2
@@ IV-C Bug Generator
11
12foreach $Response\in Responses$ do
@@ 21                   $break$;
BugFarm’s prompts consist of three parts. The first part is the natural language instruction asking LLM to generate the bugs. The second part provides contextual information about where to inject the bug, i.e., only to consider the least attended statements, $LAS$, for making bug-inducing changes. Finally, we include the entire method, including both signature and the method body, in the prompt. Such prompts are LLM-agnostic, i.e., they can be used with various existing LLMs, such as LLaMa [28], PaLM [18], Copilot [17], Alpaca [38], and ChatGPT [3]. The only consideration is to check if the prompt’s size matches the LLM context window. After the LLM’s response, Bug Generator component validates (1) if they are syntactically correct and (2) if the changes do not impact the attention of the model. The responses that do not pass these two checks will be discarded.
Algorithm 2 demonstrates BugFarm’s bug generation and validation approach. It starts by initializing the output variable $Bugs$ (Line 1) and getting the indices of LASs in $method$ (Line 2). Next, it adds an index to each statement in $method$ (Line 3). This will help us to refer to LASs in the prompt by number instead of including the whole statements, resulting in a reduction of the prompt size. This is specifically important for longer methods or statements, as the context window of LLMs is limited [29, 45]. Next, we will craft the prompt with the required context, i.e., indexed method and bug injection location, and send the prompt to an LLM (Lines 4-5)  [^1] The natural language part of our prompts is more complex than the example here. The readers can refer to our artifact to see the exact prompts [4]. Finally, once the LLM responds, we check if the generated bugs are syntactically correct (Lines 7-8) and if the changed statements are among the least attended statements by the model (Lines 9-14) to add them to the acceptable set of bugs (Lines 15-16).
V Evaluation
To evaluate the effectiveness of BugFarm, we compare it with two state-of-the-art alternative approaches, namely LEAM and $\mu$BERT, investigating the following research questions  [^2] The detailed results and additional research questions, e.g., ablation study demonstrating the necessity of prompt crafting, are publicly available [4].:
@@ 1. RQ1:
    Characteristics of the Generated Bugs. To what extent can BugFarm and alternative techniques successfully inject bugs into arbitrary code? What are the characteristics of the generated bugs by each approach?
@@ 3. RQ3:
    Effectiveness in Generating Hard-to-Repair Bugs. To what extent do learning-based Automated Program Repair (APR) techniques repair BugFarm bugs compared to those generated by alternative approaches?
@@ 4. RQ4:
V-A Experiment Setup and Data Availability
@@
Mutant Generation. We compare BugFarm with two most recent mutant generation techniques, $\mu$BERT [24] and LEAM [39]. To inject mutants, $\mu$BERT selects AST nodes representative of program behavior—literals, identifiers, expressions, assignments, object fields, method calls, array access, and types. Then, it replaces the tokens in selected AST nodes with the special token <mask> and uses CodeBERT to predict the masked token. The intuition is that if CodeBERT predicts a token different from the original one, the transformation introduces a bug. LEAM is a deep learning-based technique that learns to mutate code from large examples of real-world bugs. To that end, they represent code as a sequence of AST nodes and learn to apply eight grammar rules to select and modify the code. Both $\mu$BERT and LEAM claim to generate better bugs (mutants confirmed by test execution) compared to classic approaches, namely PIT [12] and Major [22]. Therefore, we did not include classic techniques in our evaluation.
@@
For BugFarm, we used the threshold value $k=10\%$, mainly because alternative approaches do not change more than a handful of statements in a given code (more details in section V-B). Our experimental results in the rest of this section confirm that even with such a low threshold, we still surpass other techniques in generating hard-to-detect (section V-C) and repair (section V-D) bugs. A higher threshold will make our bugs more complex compared to other approaches, improving margins. We also configured BugFarm to generate at most three mutants per method ($N=3$ in Algorithm 2) as our preliminary results show that with a higher number, the generated bugs for $k=10\%$ are almost identical. With a higher threshold value, one can also increase $N$ and generate more diverse bugs. The current implementation of BugFarm’s Bug Generator component uses GPT-3.5-turbo due to its accessibility and effectiveness at a reasonable cost (\$$\$0.0004$ per mutant on average). Using a more advanced LLM such as GPT-4 likely yields better results.
Automated Bug Confirmation. BugFarm and alternative approaches generate mutants, which entails implementing a confirmation process to ensure syntactical correctness and rule out equivalent mutants. Following the related work [39, 24], our confirmation procedure follows three steps: (1) running existing test suites on the original project and selecting the green tests; (2) compiling the generated mutants from the methods covered by the green tests and discarding syntactically incorrect ones; and (3) re-executing previously passed tests on syntactically correct mutants and choosing those killed through test execution. Due to the possibility of generating mutants that are not equivalent but survive test execution, we further performed a manual investigation on a subset of survived mutants. This will ensure the fairness of the experiments and the generalizability of the claims in the paper. All the experiments were conducted with killed mutants, which we refer to as confirmed bugs. The mutant generation techniques generated a total $1,908,566$ mutants, out of which $699,575$ were syntactically correct, and $434,215$ identified as confirmed bugs through testing (details in section V-B).
Manual Bug Confirmation. From the $265,360$ syntactically correct survived mutants, we sampled $300$, controlling equal contributions from different sets of generated mutants. We then crowd-sourced the task of mutant evaluation to $97$ individuals with varying software engineering expertise, from beginner to expert with at least two years of industry experience. We provided each participant with the mutant and original codes and asked them to label mutants with one of Yes (the mutant is a bug), No (the mutant is equivalent), or Maybe (uncertain whether the mutants is equivalent) labels. Individuals monitored their time and reported the amount spent on the task. We also asked them to evaluate how challenging it was to decide on the assigned mutants’ labels. Each mutant was labeled by three individuals, and we decided on the final label based on the majority vote.
The labeling process took $290$ person-hours and participants delivered $900$ labels ($487$ Yes, $384$ No, and $29$ Maybe). The majority voting provided us with $143$ additional confirmed bugs, where $72$ $(50.35\%)$, $53$ $(37.06\%)$, and $18$ $(12.59\%)$ of them belong to BugFarm, $\mu$BERT, and LEAM, respectively. During the labeling, we asked participants to identify the difficulty level for labeling each mutant, i.e., whether it was challenging for them to identify the mutant was equivalent or not. Figure 3 shows that BugFarm’s mutants were more challenging than other approaches for the participant to label. We believe that this large-scale human study accounts for the threat to the validity of choosing killed mutants as confirmed bugs.
Figure 3: Evaluating the complexity of generated mutants by human subjects
@@
Bug Prediction Models. When selecting bug prediction models, we had to consider the following criteria: the current implementation of BugFarm’s Attention Analyzer works on transformer-based models and requires the availability of models (weights and internals) to perform the attention analysis (we will consider including closed-source models as future work). We could not find any custom bug prediction model publicly or per request available with such characteristics. Consequently, we chose three pre-trained code-language models that are widely used by the research community for fine-tuning, namely CodeBERT [15], CodeT5 [43], and NatGen [11] [^3] Fine-tuning larger models requires non-trivial computing resources. Our experiments will show that models superior in other code-related tasks (e.g., CodeT5 over CodeBERT) show the significance of BugFarm better. We expect this to hold for larger models as well.. We fine-tuned these models using real-world bugs and evaluated the effectiveness of the outcome in predicting bugs generated by different techniques.
@@
CodeBERT is an encoder-only transformer model based on BERT [13] architecture, which is trained on $2.1M$ bi-modal (natural language and code pairs) and $6.4M$ uni-modal (code only) data from CodeSearchNet [19] dataset. The main learning objectives in CodeBERT are Masked Language Modeling (MLM)—the model learns to predict the tokens replaced by a special mask token—and Replaced Token Detection (RTD)—the model learns to detect which token does not belong to the original data. CodeT5 is an encoder-decoder transformer model based on T5 [32] architecture. CodeT5 is trained on $8.35M$ functions from various programming languages provided by CodeSearchNet [19] and BigQuery [1] dataset, and its training objectives include masked span and masked identifier prediction. Such objectives enable the model to understand code semantics better than CodeBERT.
@@
NatGen is also an encoder-decoder transformer model, trained on a generative task of naturalizing source code. Specifically, NatGen starts with CodeT5—based model—parameters and continues the training with a new objective, i.e., re-constructing the original code (natural) given transformed code (de-natural). It uses $8.1M$ pairs of natural and de-natural functions from CodeSearchNet [19] and C/C#. We used the base models of each for our experiments.
@@
Automated Program Repair Model. To evaluate the effectiveness of learning-based techniques in repairing generated bugs, we used FitRepair [44], a state-of-the-art APR technique that outperforms existing approaches. It leverages information retrieval and static analysis to implement domain-specific fine-tuning and prompting strategies. Per the authors’ instructions, we used FitRepair with CodeT5-Large in a zero-shot manner to generate the patches for $\mu$BERT, LEAM, and BugFarm confirmed bugs and validated patches through test execution.
@@
Subjects. BugFarm is programming-language agnostic; none of its components depend on a specific programming language. However, we chose Java projects in our experiments for the following reasons: (1) LEAM’s pre-trained model is on Java, and they have no alternative training dataset for other programming languages; (2) we needed real-world bug datasets for RQ2, and most of the existing real-world bug datasets are in Java. For a fair comparison, we used Defects4J $V2.0$ projects (but used their latest version) as a baseline for bug generation since the alternative approaches are shown to work on them with no issues. The current version of BugFarm supports Maven projects only, so we excluded Mockito and Closure projects from the subjects. The first two columns of Table I show the list of our $15$ subjects and the number of methods per subject used for mutant generation.
@@
TABLE I: Comparing the characteristics of bugs generated by BugFarm, $\mu$BERT, and LEAM. M: # Methods, SCM: # Syntactically Correct Mutants, CB: # Confirmed Bugs, SI: # Statements Involved in bug generation, LD: Lines Deleted in bug generation, ED: Edit Distance, OL: Overlap with LEAM, OM: Overlap with $\mu$BERT.
Subjects M BugFarm LEAM [39] $\mu$BERT [24]
SCM CB SI LD ED OL OM SCM CB SI LD ED SCM CB SI LD ED
cli 276 295 266 2.9 0.0% 33.2 $\langle$13.9%,0.6$\rangle$ $\langle$8.6%,0.7$\rangle$ 2126 1388 2.6 13.69% 26.1 2053 1532 2.0 0.07% 19.3
codec 901 544 360 3.1 0.0% 31.6 $\langle$3.1%,0.7$\rangle$ $\langle$5.5%,0.9$\rangle$ 4607 2130 2.8 7.84% 22.4 14738 9551 2.0 0.0% 17.9
collections 4440 3775 3175 2.9 0.0% 29.7 $\langle$11.1%,0.6$\rangle$ $\langle$5.1%,0.7$\rangle$ 21379 11266 2.8 7.77% 20.9 34038 22196 2.0 0.0% 20.0
compress 4123 573 468 3.0 0.0% 29.4 $\langle$4.4%,0.7$\rangle$ $\langle$2.4%,0.8$\rangle$ 20401 9373 2.6 12.55% 24.5 57815 23602 2.0 0.0% 24.2
csv 248 282 252 2.7 0.0% 24.9 $\langle$11.9%,0.7$\rangle$ $\langle$1.5%,0.7$\rangle$ 1803 1189 2.8 14.55% 33.5 131 109 2.0 0.0% 7.0
jxpath 1672 1152 948 3.3 0.0% 28.0 $\langle$8.0%,0.7$\rangle$ $\langle$6.1%,0.8$\rangle$ 11511 6169 2.9 15.24% 26.2 21335 10907 2.0 0.0% 25.8
lang 3810 4421 3791 3.0 0.0% 30.1 $\langle$7.5%,0.7$\rangle$ $\langle$3.1%,0.8$\rangle$ 26365 16316 2.7 9.69% 21.4 23312 15732 2.0 0.02% 21.1
math 5796 6466 6207 3.4 0.0% 34.2 $\langle$4.4%,0.7$\rangle$ $\langle$2.3%,0.8$\rangle$ 43090 42262 2.7 8.43% 17.0 85393 83512 2.0 0.01% 18.3
gson 985 631 545 3.3 0.0% 38.2 $\langle$7.3%,0.7$\rangle$ $\langle$1.6%,0.7$\rangle$ 4896 2992 2.6 17.95% 27.9 4173 2067 2.0 0.05% 19.8
jackson-core 2626 2281 1732 4.0 0.0% 43.1 $\langle$2.9%,0.7$\rangle$ $\langle$2.3%,0.7$\rangle$ 21472 9963 2.5 22.58% 25.8 19118 9310 2.1 0.0% 21.6
jackson-db 8076 4828 3965 3.9 0.0% 41.9 $\langle$7.2%,0.7$\rangle$ $\langle$6.9%,0.7$\rangle$ 36155 18454 2.8 12.17% 28.7 51032 22454 2.0 0.0% 22.3
jackson-xml 586 312 258 3.9 0.0% 44.0 $\langle$6.3%,0.7$\rangle$ $\langle$5.9%,0.7$\rangle$ 3096 1580 2.7 12.85% 31.1 2576 1226 2.0 0.08% 14.7
jfreechart 8602 4051 3186 3.1 0.0% 27.1 $\langle$7.4%,0.7$\rangle$ $\langle$1.1%,0.8$\rangle$ 40313 17987 2.5 10.64% 19.8 18924 7045 2.0 0.07% 28.2
joda-time 4279 4188 3734 2.6 0.0% 22.1 $\langle$8.6%,0.6$\rangle$ $\langle$3.6%,0.7$\rangle$ 26224 15641 2.9 6.29% 24.7 49907 28694 2.0 0.0% 21.1
jsoup 1642 1561 1356 2.8 0.0% 28.5 $\langle$9.4%,0.7$\rangle$ $\langle$4.7%,0.7$\rangle$ 9534 5456 2.5 14.15% 23.6 6699 4013 2.2 0.05% 21.0
Total 48062 35359 30242 - - - - - 272972 162166 - - - 391244 241950 - - -
Average 3204 2357 2016 3.2 0.0% 32.4 $\langle$7.6%,0.7$\rangle$ $\langle$4.0%,0.7$\rangle$ 18198 10811 2.7 12.43% 24.9 26083 16130 2.0 0.02% 20.2
@@
Evaluation Metrics. To compare the performance of bug prediction models, we use accuracy, precision, recall, F1 score, and False Negative Rate (FNR) as our metrics. To evaluate the APR results, we measure the repair success rate, i.e., the number of bugs the technique successfully patches. True Positive (TP) is when the code is buggy, and the model predicts it as buggy. True Negative (TN) is when the code is not buggy, and the model predicts it as non-buggy. False Positive (FP) is when code is not buggy, but the model predicts it as buggy. False Negative (FN) is when the code is buggy, but the model predicts it as non-buggy.
V-B RQ1: Characteristics of the Generated Bugs
Table I presents the metrics and their values calculated for the studied approaches (For BugFarm, the reported number is aggregated or averaged across all baseline models). Columns SCM and CB show the number of syntactically correct mutants and confirmed bugs, respectively. The mutant generation techniques overall generated a total of $1,908,566$ mutants, out of which $699,575$ were syntactically correct and $434,358$ confirmed as bugs (through test execution and human study). The percentage of syntactically correct to all generated mutants in BugFarm is $85\%$ compared to that of $59.4\%$ and $61.82\%$ for LEAM and $\mu$BERT, corroborating a higher quality of our mutants. A higher percentage of confirmed bugs ($85.55\%$ in BugFarm compared to $59.41\%$ and $61.84\%$ for LEAM and $\mu$BERT), along with the human study results presented in section V-A, show the code transformations by BugFarm are more likely to be bugs. The rest of the metrics are computed for confirmed bugs, which we simply refer to as bugs in the rest of the paper: